Suppose that Paulie and Vinny each can produce ice cream or t-shirts. The table shows the quantity of each good that Paulie and Vinny each can produce in 11 hour, respectively, if they devote all of their time and effort into making the good. Round all answers to two decimal places.

Ice cream (cups) T-shirts (quantity)
Paulie 3.003.00 20.0020.00
Vinny 7.007.00 14.0014.00

a. What is Paulie's opportunity cost of producing a cup of ice cream?

Paulie's opportunity cost:

t‑shirtst‑shirts

b. What is Vinny's opportunity cost of producing a t-shirt?

Vinny's opportunity cost:

cups of ice creamcups of ice cream

c. Paulie has a comparative advantage in

and Vinny has a comparative advantage in
Answer #1

(a)

Paulie's opportunity cost of a cup of ice-cream = 20.00 / 3.00 = 6.67 T-shirts

(b)

Vinnie's opportunity cost of a T-shirt = 7.00 / 14.00 = 0.50 cups of ice cream

(c)

Paulie's opportunity cost of a T-shirt = 3.00 / 20.00 = 0.15 cups of ice cream

Therefore,

Paulie has comparative advantage in T-shirt (since she can make T-shirts at lower opportunity cost than Vinnie can: 0.15 < 0.5).

Vinnie has comparative advantage in ice creams.